OpenStudy (grant330sims):

Explain, in complete sentences, how you would use the graphing method to solve the following system of inequalities and what the final solution will look like on the coordinate plane. x + y –6 2x – 3y < 0

OpenStudy (aroub):

What's the sign for the first equation?

OpenStudy (grant330sims):

my bad. its a greater than or equal to.

OpenStudy (grant330sims):

is it really that hard.... And when i say that i dont mean to be rude.. just kinda shocked

OpenStudy (aroub):

To graph a linear inequality in two variables the first thing you got to do is convert each inequality to it's associated equation ( for every inequality has an associated equation ) lets do one by one: x+y>=-6 => x+y=-6  => (0,-6) and (-6,0) Since the inequality is not strict, the line is drawn solid.  (0,0) is in the half-plane above the line. Substitute its coordinates in the inequality: 0+0>=-6 is true. Since the coordinates of (0,0) does satisfy the linear inequality, you shade the other half-plane that does not contain the points (0,0)